OpenStudy (kweenlexis):

The last question was the wrong question the the one i need help in PLEASE HELP ME. Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5)

OpenStudy (youngstudier):

What do you think the answer is @kweenlexis ?

OpenStudy (kweenlexis):

A??

OpenStudy (youngstudier):

Um, no, but you're close.

OpenStudy (kweenlexis):

D??

OpenStudy (youngstudier):

No, you're getting closer though. The way you group or Associate the numbers does not change the result.

OpenStudy (kweenlexis):

its it B?!

OpenStudy (youngstudier):

No, B is an identity property.

OpenStudy (youngstudier):

Associative property of Addition: (a + b) + c = a + ( b+ c)

OpenStudy (kweenlexis):

oooooh now i see.

OpenStudy (youngstudier):

So the answer would be C, since it's the only answer choice left. :)

OpenStudy (kweenlexis):

correct or incorrect? If i had (6+2)+3=6+(2+3)

OpenStudy (youngstudier):

That would be correct. :)

OpenStudy (kweenlexis):

ok

OpenStudy (youngstudier):

You'd still get the same results :)

OpenStudy (kweenlexis):

Thank you!!!!!!!!!!!!!!!

OpenStudy (youngstudier):

You're welcome!!! :)

OpenStudy (kweenlexis):

What is the value of the expression? 12.4 + (–2 + 1.6) A. 11 B. 12 C. 15 D. 16

OpenStudy (kweenlexis):

isn't B the answer.

OpenStudy (youngstudier):

One sec.

OpenStudy (kweenlexis):

ok

OpenStudy (youngstudier):

Yuppers, B is correct :) Good job

OpenStudy (kweenlexis):

yay about time i got it right..thank you.

OpenStudy (kweenlexis):

im going to stop asking you question now so your head wont start hurting and stuff.

OpenStudy (youngstudier):

You're welcome, you did that one by yourself, I just double-checked the answer :)

OpenStudy (kweenlexis):

yep. 2 points for you 1 for me lol.

OpenStudy (youngstudier):

Points?

OpenStudy (kweenlexis):

nothing but can i ask you 1 more i promise this wioll be the last 1.

OpenStudy (youngstudier):

Sure! :D

OpenStudy (kweenlexis):

will*

OpenStudy (kweenlexis):

What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17 I got c.

OpenStudy (youngstudier):

C would be correct :)

OpenStudy (kweenlexis):

yay!

OpenStudy (youngstudier):

:) Good job!

OpenStudy (kweenlexis):

ty

OpenStudy (youngstudier):

You're welcome!

OpenStudy (kweenlexis):

alrighty

OpenStudy (kweenlexis):

I MADE A 100% thank you soooo much!